# Constrained Language Generation with Discrete Diffusion Models

### v1

@@ 5.3.2 Novelty Constraint
As shown in Table 2, the CBG setting yields a 203.4% increase in novel molecule generation, while the CFG setting shows a 161.4% increase. Even without guidance, the method still produces 48.1% more novel molecules, with only a minimal reduction in the QED score.
These results demonstrate CDD’s applicability for scientific applications. Ensuring novelty is imperative for molecule discovery, while chemical property adherence is crucial for real world applications in drug-design or materials discovery.
6 Limitations
@@
7 Conclusion
In this paper we presented Constrained Discrete Diffusion (CDD), a novel method for imposing both structured and unstructured constraints during language generation. CDD relies on the adoption of discrete diffusion models which enable the consideration of entire sequences when imposing constraints during the generation process. Offering a key departure from current paradigms for controllable text generation, CDD integrates differentiable optimization methods into the diffusion reverse process. This integration, governed by a rigorous mathematical optimization framework, provides precise control for natural language generation. The method is empirically validated on (i) toxicity mitigation and harmful content prevention, (ii) imposition of lexical constraints at character and sequence levels, and (iii) generation of novel molecular sequences that adhere to specific properties for drug discovery, reporting state-of-the-art results for controllable generation in both terms of constraint adherence and quality metrics.
Acknowledgments
Appendix A Expanded Related Work
@@
Appendix B Experimental Setup
B.1 Toxicity Mitigation
@@
As the surrogate model for toxicity task, we use a GPT-Neo (1.3B) model, adapted for binary classification. We finetune this model on the Jigsaw toxicity dataset which includes multiple toxicity-related labels such as toxic, severe toxic, insult, etc. We consolidates these columns into a single binary target (toxic vs. non-toxic).
B.2 Counting Constraints
Instruction Tuning Dataset.
@@ B.2 Counting Constraints
B.3 Lexical Constraints
Instruction Tuning Dataset.

### v2

@@ 5.2 Molecular Generation
We compare agasint baselines with no QED guidance alongside CFG and CBG QED guidance [8] to evaluate the impact of our novelty constraint. The results are shown in the right side of Figure 5. Notice how the CBG setting yields a 203.4% increase in novel molecule generation, while the CFG setting shows a 161.4% increase. Even without guidance, the method still produces 48.1% more novel molecules, with only a minimal reduction in the QED score. These results are significant because they demonstrate that CDD can effectively generate novel molecules while maintaining a high QED score, which is crucial for many scientific tasks.
5.3 Instruction Following
Model Size PPL LLM–Judge Viol (%)
Counting GPT-2 (zero-shot) 124M 19.2 ± 0.8 57.8 100
GPT-2 (instruction-tuned) 124M 23.4 ± 1.3 70.9 5.0
Llama 3.2 (zero-shot) 1B 34.3 ± 3.8 79.5 100
Llama 3.2 (instruction-tuned) 1B 30.7 ± 2.9 91.8 9.1
MDLM 169M 16.9 ± 1.0 61.2 54.5
CDD (Ours) 169M 17.9 ± 1.1 60.4 0.0
Lexical GPT-2 (zero-shot) 124M 26.6 ± 4.2 28.1 97.6
GPT-2 (instruction-tuned) 124M 62.2 ± 2.9 15.4 96.4
Llama 3.2 (zero-shot) 1B 11.3 ± 2.0 30.7 97.2
Llama 3.2 (instruction-tuned) 1B 66.5 ± 6.1 22.9 93.6
MDLM 169M 21.9 ± 2.0 34.8 97.5
CDD (Ours) 169M 26.6 ± 4.1 65.1 0.0
Table 1: PPL is assessed using GPT2. Bold and underlined mark the best and second-best, respectively.
Finally, the last application focuses on satisfying user-specified symbolic constraints, a task that many autoregressive models struggle with. We evaluate CCD against such models using two types of constraints. First, we assess the ability of the tested models to accurately answer how many letters are in a word, such as ‘How many R’s are in Strawberry?’. We additionally test lexical constraints based of the COLLIE dataset [53], prompting the models to generate word-indexed constrained sentences, such as ‘Generate a sentence where the third-to-last word is mankind’. For each constraint, the projection operator to imposes hard constraints on the generation, pushing the samples back onto the constraint set.
Counting constraints. As illustrated in the top of Table 1, unconstrained GPT-2 and Llama 3.2 fail to satisfy any letter-count requirement. Instruction tuning helps, reducing violation rates to 5.0% (GPT-2) and 9.1% (Llama 3.2), but still leaves a notable error margin. Alternatively, MDLM achieves the lowest perplexity and second-best coherence, but over half of its outputs violate the constraint. By contrast, CDD fully eliminates violations while preserving near-best fluency (17.9 perplexity), demonstrating its provision of strict compliance without sacrificing language quality.
Lexical constraints. The results for lexical constraints are shown in the bottom of Table 1. The zero-shot and instruction-tuned versions of GPT-2 and Llama 3.2 show similarly high violation rates (about 97%), and instruction tuning only minimally improves these figures. MDLM again delivers relatively high fluency (21.9 perplexity, 34.8 coherence) but fails to address the precise positional requirements (97.5% violations). In stark contrast, CDD ensures perfect adherence and the best LLM-as-a-Judge score (65.1 coherence), with only a moderate increase in perplexity, thus confirming the effectiveness of its projection mechanism for enforcing hard lexical constraints. Taken together, these experiments reveal a fundamental limitation of conventional and instruction-tuned language models in scenarios requiring strict constraints. Conversely, discrete diffusion models, when equipped with constraints, can effectively produces outputs that satisfy both counting and lexical requirements, opening new avenues for applications in controlled text generation.
We include more evaluation and experiments in the Appendix. Specifically, Appendix C details the experimental setup, Appendix D evaluates CDD’s robustness through an extensive hyper-parameter ablation, and Appendix E reports runtime comparisons with the baselines.
6 Conclusion
In this paper we presented Constrained Discrete Diffusion (CDD), a novel method for imposing both structured and unstructured constraints during sequence generation. CDD relies on the adoption of discrete diffusion models which enable the consideration of entire sequences when imposing constraints during the generation process. Offering a key departure from current paradigms for controllable text generation, CDD integrates differentiable optimization methods into the diffusion reverse process. This integration, governed by a rigorous mathematical optimization framework, provides precise control for discrete sequence generation. The method is empirically validated on (i) toxicity mitigation and harmful content prevention, (ii) generation of novel molecular sequences that adhere to specific properties for drug discovery, and (iii) imposition of lexical constraints at character and sequence levels, reporting state-of-the-art results for controllable generation in both terms of constraint adherence and quality metrics.
Acknowledgments
Appendix A Limitations
@@
Appendix B Expanded Related Work
@@
Appendix C Experimental Setup
C.1 Toxicity Mitigation
LLM-as-a-Judge Setup.
To quantify the fluency and coherence of generated continuations we follow the increasingly adopted LLM-as-a-Judge evaluation. This method has been shown to result in similar evaluation as human judgment [64]. Specifically, each post-prefix sentence is scored by the instruction-tuned GEMMA-3-27B-IT model and issued the following prompt:
“You are a language expert evaluating the fluency and coherence of the following AI-generated sentence… Give a single integer score between 0 (poor) and 100 (excellent), with no explanation or comments.”
The model returns a single scalar $s\in[0,100]$ for each sentence. Using an LLM evaluator offers two key advantages: (i) strong correlation with human judgments at negligible annotation cost, and (ii) scalability for thousands of evaluations performed in various experiments.
@@ C.1 Toxicity Mitigation
As the surrogate model for toxicity task, we use a GPT-Neo (1.3B) model, adapted for binary classification. We finetune this model on the Jigsaw toxicity dataset which includes multiple toxicity-related labels such as toxic, severe toxic, insult, etc. We consolidate these columns into a single binary target (toxic vs. non-toxic).
PPLM Implementation.
As a baseline for toxicity reduction in natural language generation we include the PPLM methodology [26] . PPLM uses a frozen pretrained GPT-2 generator for output steering. At every decoding step the algorithm runs the usual forward pass, takes a few gradient-ascent steps on the cached key/value activations to maximize the score of a small attribute model which in this case is a toxicity classifier trained on the Jigsaw toxicity dataset. It simultaneously keeps the perturbed logits close to the original distribution with a KL penalty and a geometric-mean fusion term. We use the authors implementation and code structure without architectural modifications. The underlying steering mechanism, loss terms and hyper-parameter values are unchanged from the authors’ baseline.
FUDGE Implementation.
For another toxicity control baseline we additionally implemented FUDGE [15] on the Jigsaw Toxicity dataset. Following the authors implementation we use an LSTM followed by a linear output layer trained to predict future toxicity from partial prefixes. At inference time we add its scaled log non-toxicity probability to GPT-2’s next-token logits for the top candidates then sample from the re-weighted distribution. While the original implementation of FUDGE fixes the guidance strength $\lambda=1$, we perform a hyperparameter seach and sample from the resulting distribution with $\lambda\in\{1,\dots,9\}$, displaying the toxicity and perplexity results in Table 2. We use the authors’ original code structure, and base language model.
$\lambda$ PPL (Avg) Coherence Viol % $\tau=0.25$ Viol % $\tau=0.5$ Viol % $\tau=0.75$
1 20.34 41.64 33.5 20.0 13.1
2 26.46 41.79 31.6 19.7 12.7
3 36.36 40.07 33.1 20.7 13.3
4 43.80 40.37 31.8 21.5 13.3
5 54.37 40.12 31.7 21.6 13.0
6 65.23 39.33 33.1 21.7 13.1
7 75.23 38.76 32.5 20.3 12.9
8 76.51 39.09 32.2 20.8 12.6
9 81.84 38.91 30.6 19.6 11.4
Table 2: FUDGE perplexity, coherence, and violation rates at three thresholds for varying $\lambda$.
Diffusion Guidance Implementation.
While recent literature [8] has proposed adaptations of classifier-based guidance [28] and classifier-free guidance [65] for discrete diffusion models, these approaches have only been employed for settings with limited vocabulary sizes (e.g., molecule generation where the vocabulary is composed of 40 tokens). For natural language tasks which have much larger vocabulary sizes, for instance, the GPT2 tokenizer is composed of over 50,000 tokens–these techniques have not been applied. Currently, this is a scalability issue, as these approaches require that guidance vectors are computed of size $\mathbb{R}^{L\times N}$, making it much simpler and efficient when operating on a limited vocabulary size. Hence, as there is no prior literature to the extent of our knowledge that has scaled these techniques to natural language tasks, we defer comparison to these methods for the molecular generation setting.
Implementation Details.
For the ALM projection, we initialize using the following hyperparameters: $\lambda_{\text{init}}=0.0,\mu_{\text{init}}=1.0,\mathrm{outer\_iter}_{\text{max}}=1000,\mathrm{inner\_iter}_{\text{max}}=10,\eta=0.20$. We choose these setting for the toxicity experiment as they result in the lowest perplexity in the Appendix D ablation study. For natural language toxicity reduction details we use a generation length of 100 tokens.
C.2 Molecular Generation
Training Dataset.
C.2.1 Synthetic Accessibility Constraints
Surrogate Model.
FUDGE Implementation.
For this setting, we follow the FUDGE implementation provided by [8]. However, while their guidance is trained on QED scores, labeling samples from the QM9 dataset with these scores, we adapt it to label the training data with RDKIT’s computation of the synthetic accessibility scores.
Diffusion Guidance Implementation.
Similar to the implementation of FUDGE, we adapt the guidance mechanisms provided by [8] for QED by retraining on synthetic accessibility labels. Otherwise, our implementation mirrors their approach for QED guidance.
ALM Implmenetation.
For the ALM projection, we initialize using the following hyperparameters: $\lambda_{\text{init}}=0.0,\mu_{\text{init}}=1.0,\mu_{\text{max}}=1000,\mathrm{outer\_iter}_{\text{max}}=1000,\mathrm{inner\_iter}_{\text{max}}=100,\eta=1.0$.
C.2.2 Novelty Constraints
Novelty Projection Operator.
FUDGE Implementation.
For this application, we similarly follow the FUDGE implementation and configuration in [8] for QED guidance.
Diffusion Guidance Implementation.
We use classifier based guidance and classifier free guidance for QED as implemented in [8].
C.3 Instruction Following
LLM-as-a-Judge Setup.
To quantify the fluency and coherence of generated continuations we follow the increasingly adopted LLM-as-a-Judge evaluation. This method has been shown to result in similar evaluation as human judgment [64]. Specifically, each post-prefix sentence is scored by the instruction-tuned Gemini-2.0-Flash model and issued the following prompt:
“You are a language expert evaluating the fluency and coherence of the following AI-generated sentence… Give a single integer score between 0 (poor) and 100 (excellent), with no explanation or comments.”
The model returns a single scalar $s\in[0,100]$ for each sentence. Using an LLM evaluator offers two key advantages: (i) strong correlation with human judgments at negligible annotation cost, and (ii) scalability for thousands of evlautions performed in various experiments.
Discrete Diffusion Base Model.
While in Section 5.1, we directly utilize the pretrained MDLM checkpoint provided by [7], for the experiments in Section 5.3, we pretrain and subsequently instruction-tune a slightly larger MDLM model, adopting the Llama 1 tokenizer. Empirical observations indicated that employing a marginally larger model and vocabulary size was essential for optimal performance on instruction-guided tasks. Consequently, we adopt this enhanced model for all subsequent evaluations in this context, including both the base MDLM experiments and the proposed CDD framework. Our instruction tuning is conducted using the ShareGPT dataset, and all pretraining follows the design choices and hyperparameters provided by [7].
C.3.1 Counting Constraints
Instruction-Tuning Dataset.
@@ C.3.1 Counting Constraints
C.3.2 Lexical Constraints
Instruction-Tuning Dataset.